OpenStudy (anonymous):

What is the surface area of the square pyramid? A. 2242 in2 B. 3838 in2 C. 4636 in2 D. 30,324 in2

OpenStudy (anonymous):

http://prntscr.com/9q5pw6

OpenStudy (anonymous):

@purple_pink

jabez177 (jabez177):

Find the surface area with a base length of 3 cm, and a slant height of 2 cm SA = s2 + 2 × s × l SA = 32 + 2 × 3 × 2 SA = 9 + 12 SA = 21 cm2

jabez177 (jabez177):

Read example #2, it really helps you learn. :)

OpenStudy (purple_pink):

k i'm here

OpenStudy (anonymous):

hi

OpenStudy (purple_pink):

lemme calculate first, give me 2 minutes

OpenStudy (anonymous):

k

OpenStudy (trevor12):

littleleefie im your fan

OpenStudy (anonymous):

ok ty

OpenStudy (trevor12):

welcome do you know how to private chat

OpenStudy (anonymous):

no

OpenStudy (trevor12):

see the little mail box

OpenStudy (anonymous):

oh then yes i do know how to private chat

OpenStudy (trevor12):

so did you figure it out now

OpenStudy (anonymous):

the question

OpenStudy (purple_pink):

i think C but i'm not so sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have 1 more

OpenStudy (purple_pink):

k